40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sun Aug 5 13:28:43 2018, gautam, Update, CDS, c1lsc flaky 
    Reply  Mon Aug 6 00:26:21 2018, gautam, Update, CDS, More CDS woes CDSoverview.png
       Reply  Mon Aug 6 14:38:38 2018, gautam, Update, CDS, More CDS woes CDScrash.pngMC1failures.png
          Reply  Mon Aug 6 19:49:09 2018, gautam, Update, CDS, More CDS woes IMG_7106.JPG
             Reply  Tue Aug 7 11:30:46 2018, gautam, Update, CDS, More CDS woes 
                Reply  Tue Aug 7 22:28:23 2018, gautam, Update, CDS, More CDS woes 
                   Reply  Wed Aug 8 23:03:42 2018, gautam, Update, CDS, c1lsc model started 
                      Reply  Thu Aug 9 12:31:13 2018, gautam, Update, CDS, CDS status update 
                         Reply  Wed Aug 15 21:27:47 2018, gautam, Update, CDS, CDS status update 
                            Reply  Tue Sep 4 10:14:11 2018, gautam, Update, CDS, CDS status update 
                               Reply  Wed Sep 5 10:59:23 2018, wgautam, Update, CDS, CDS status update 
                                  Reply  Thu Sep 6 14:21:26 2018, gautam, Update, CDS, ADC replacement in c1lsc expansion chassis CDSoverview.png
                                     Reply  Fri Sep 7 12:35:14 2018, gautam, Update, CDS, ADC replacement in c1lsc expansion chassis Screenshot_from_2018-09-07_12-34-52.png
                                        Reply  Mon Sep 10 12:44:48 2018, Jon, Update, CDS, ADC replacement in c1lsc expansion chassis 
                                     Reply  Thu Sep 20 11:29:04 2018, gautam, Update, CDS, New PCIe fiber housed PCIeFiberSwap.pngPCIeFiberSwap_FBrebooted.png
                                        Reply  Thu Sep 20 16:19:04 2018, gautam, Update, CDS, New PCIe fiber install postponed to tomorrow 
                                           Reply  Fri Sep 21 16:46:38 2018, gautam, Update, CDS, New PCIe fiber installed and routed PCIeFiber.pngIMG_5878.JPG
Message ID: 14206     Entry time: Fri Sep 21 16:46:38 2018     In reply to: 14203
Author: gautam 
Type: Update 
Category: CDS 
Subject: New PCIe fiber installed and routed 

[steve, koji, gautam]

We took another pass at this today, and it seems to have worked - see Attachment #1. I'm leaving CDS in this configuration so that we can investigate stability. IMC could be locked. However, due to the vacuum slow machine having failed, we are going to leave the PSL shutter closed over the weekend.

Attachment 1: PCIeFiber.png  22 kB  Uploaded Fri Sep 21 17:47:07 2018  | Hide | Hide all
PCIeFiber.png
Attachment 2: IMG_5878.JPG  1.707 MB  Uploaded Fri Sep 21 17:55:51 2018  | Hide | Hide all
IMG_5878.JPG
ELOG V3.1.3-